At Wendy’s there are 11 options for burger toppings. Assuming you want a single burger and cannot...

  1. At Wendy’s there are 11 options for burger toppings. Assuming you want a single burger and cannot do double toppings, how many different burgers can you order that have
    a. 0 toppings
    b.1 topping
    c. 2 toppings
    d. 6 toppings
    e. 8 toppings
    f. 9 toppings
    g. 10 toppings
    h. 11 toppings
Answer #1

Number of burgers each with k toppings out of 11 are  11Ck

a.

Number of burgers each with 0 toppings = 11C0   = 11! / ((11 - 0)! * 0!) = 1

b.

Number of burgers each with 1 topping1 = 11C1   = 11! / ((11 - 1)! * 1!) = 11! / 10! = 11

c.

Number of burgers each with 2 toppings = 11C2   = 11! / ((11 - 2)! * 2!) = (11 * 10) / 2 = 55

d.

Number of burgers each with 6 toppings = 11C6   = 11! / ((11 - 6)! * 6!) = (11 * 10 * 9 * 8 * 7) / (5 * 4 * 3 * 2 * 1) = 462

e.

Number of burgers each with 8 toppings = 11C8   = 11! / ((11 - 8)! * 8!) = (11 * 10 * 9) / (3 * 2 * 1) = 165

f.

Number of burgers each with 9 toppings = 11C9   = 11! / ((11 - 9)! * 9!) = (11 * 10) / ( 2 * 1) = 55

g.

Number of burgers each with 10 toppings = 11C0   = 11! / ((11 - 10)! * 10!) = 11 / 1 = 11

h.

Number of burgers each with 11 toppings = 11C11   = 11! / ((11 - 11)! * 11!) = 1